# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@ 3.3. Geodesic convexity in the Fisher-Rao geometry
Finally, to smooth $\rho$ in (3.22), we introduce the mollifier:
@@
which is smooth, has a compact support within $B_{\delta}$, and satisfies $\int\phi_{\delta}(\theta)\mathrm{d}\theta=1$. The smoothed density is constructed using convolution as follows:
$\rho_{\delta}(\theta)={\rho^{*}}(\theta)\bigl{(}\phi_{\delta}*(x_{1}\mathbbm{1}_{S^{c}_{r+\delta}}+x_{2}\mathbbm{1}_{S_{r-\delta}}+x_{3}\mathbbm{1}_{S_{r+\delta}\backslash S_{r-\delta}})\bigr{)}(\theta).$
Here, $x_{3}$ is determined by $\int\rho_{\delta}(\theta)\mathrm{d}\theta=1$, and it satisfies $x_{2}\leq x_{3}\leq x_{1}$, since
$\begin{split}\phi_{\delta}*(x_{1}\mathbbm{1}_{S^{c}_{r+\delta}}+x_{2}\mathbbm{1}_{S_{r-\delta}}+x_{2}\mathbbm{1}_{S_{r+\delta}\backslash S_{r-\delta}})\leq&x_{1}\mathbbm{1}_{S^{c}_{r}}+x_{2}\mathbbm{1}_{S_{r}}\\ \leq&\phi_{\delta}*(x_{1}\mathbbm{1}_{S^{c}_{r+\delta}}+x_{2}\mathbbm{1}_{S_{r-\delta}}+x_{1}\mathbbm{1}_{S_{r+\delta}\backslash S_{r-\delta}}).\end{split}$
This also leads to that $x_{2}\leq\frac{\rho_{\delta}}{{\rho^{*}}}\leq x_{1}$. Since $\rho_{\delta}(\theta)$ only deviates from $\rho$ (as denoted in (3.22)) within the compact band $S_{r+2\delta}\backslash S_{r-2\delta}$, with the measure depending on $\delta$, the estimations in eq. 3.24 remain valid when $\delta$ is sufficiently small. ∎
Theorem 3.3 demonstrates that the KL divergence, a specific instance of an $f$-divergence, is not geodesically convex. This implies that, in general, $f$-divergences do not necessarily attain such convexity. In what follows, we establish a sufficient condition on $f$ that guarantees the geodesic convexity of the corresponding $f$-divergence.
@@
Assume $f\in C^{2}(0,+\infty)$ is convex, satisfying $f(1)=0$ and $xf^{\prime}(x)$ is concave. Then, for any smooth, positive target density $\rho^{*}$, the $f$-divergence is geodesically convex within the Fisher-Rao geometry. Furthermore, when $f^{\prime\prime}(1)>0$, such an $f$-divergence is $\alpha_{f}$-strongly geodesically convex, where $\alpha_{f}>0$ is a constant that depends solely on $f$. These assumptions are satisfied by $f^{\prime\prime}(x)=x^{p}$ for $p\leq-2$.
Proof.
@@
(3.25) $\mathbb{E}_{\rho}[f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}]=\mathbb{E}_{{\rho^{*}}}\Bigl{[}\frac{\rho}{{\rho^{*}}}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\Bigr{]}\leq\mathbb{E}_{{\rho^{*}}}\bigl{[}\frac{\rho}{{\rho^{*}}}\bigr{]}f^{\prime}\Bigl{(}\mathbb{E}_{{\rho^{*}}}\bigl{[}\frac{\rho}{{\rho^{*}}}\bigr{]}\Bigr{)}=f^{\prime}(1).$
We recall the expression of Hessian operator from (3.17b)
[…]
where in the last line we used (3.25). We introduce the function
(3.27) $h(x)=2xf^{\prime\prime}(x)+f^{\prime}(x)-f^{\prime}(1)=2(xf^{\prime}(x))^{\prime}-f^{\prime}(x)-f^{\prime}(1).$
@@
(3.28) $g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma)\geq\frac{1}{2}\int\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}h\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\mathrm{d}\theta$
Due to our assumptions, $xf^{\prime}(x)$ and $-f(x)$ are both concave, so $(xf^{\prime}(x))^{\prime}$ and $-f^{\prime}(x)$ are non-increasing functions of $x$. It follows that $h(x)$ (see (3.27)) is non-increasing. Therefore, when $x\leq 1$, we have
$h(x)\geq h(1)=2f^{\prime\prime}(1)\geq 0,$
and when $x\geq 1$, we have
$h(x)\geq f^{\prime}(x)-f^{\prime}(1)\geq 0.$
Therefore, $h(x)\geq 0$. This together with (3.28) implies that
$g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma)\geq 0,$
for any tangent vector $\sigma$. Thus, the $f$-divergence is geodesically convex.
Next we prove that when $f^{\prime\prime}(1)>0$, the $f$-divergence is $\alpha_{f}$-strongly geodesically convex. To do so, we need a stronger estimate on the function $h(x)$. When $f^{\prime\prime}(1)>0$, there exists $\delta_{f}>0$, such that for $x\in[1,1+\delta_{f}]$,
$\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x).$
This estimate, together with the discussions above, implies that
(3.29) $h(x)\geq\begin{cases}2f^{\prime\prime}(1)&0<x\leq 1\\ 2xf^{\prime\prime}(x)\geq f^{\prime\prime}(1)&1<x\leq 1+\delta_{f}\\ f^{\prime}(x)-f^{\prime}(1)=\int_{1}^{x}f^{\prime\prime}(\xi){\rm d}\xi\geq\delta_{f}\frac{f^{\prime\prime}(1)}{2}&x\geq 1+\delta_{f}\end{cases}.$
Therefore, $h(x)\geq\alpha_{f}>0$, where $\alpha_{f}=\min\{f^{\prime\prime}(1),\delta_{f}\frac{f^{\prime\prime}(1)}{2}\}$ is a constant that depends only on $f$. Therefore, when $f^{\prime\prime}(1)>0$, combining (3.29) and (3.28), we conclude that the $f$-divergence is $\alpha_{f}$-strongly geodesically convex, i.e.,
@@
Finally, for $f^{\prime\prime}(x)=x^{p}$ with $p\leq-2$, we have $f$ is convex on $(0,+\infty)$. Moreover, $xf^{\prime}(x)$ is concave, since $(xf^{\prime}(x))^{\prime\prime}=2f^{\prime\prime}(x)+xf^{\prime\prime\prime}(x)=(2+p)x^{p}\leq 0$ for $x\in(0,+\infty)$. Thus, all the assumptions are satisfied for $f$ with $f^{\prime\prime}(x)=x^{p}$ with $p\leq-2$. ∎
4. Functional Inequality: Gradient Dominance Condition
In this section, we turn to the gradient dominance condition (1.11) of the $f$-divergences in the Fisher-Rao geometry. Counterexamples will be presented to illustrate that the KL-divergence does not satisfy the gradient dominance condition. Then sufficient and necessary conditions will be provided to identify the $f$ that ensures the gradient dominance condition.
@@ 4. Functional Inequality: Gradient Dominance Condition
Thus, the gradient dominance condition can be formulated as
@@
There exists $\alpha_{f}>0$ depending only on $f$, such that
(4.1) $\int\rho\left(f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-\mathbb{E}_{\rho}\Bigl{[}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\Bigr{]}\right)^{2}\mathrm{d}\theta\geq\alpha_{f}D_{f}[\rho\|{\rho^{*}}].$
The above inequality holds for any smooth positive densities $\rho$ and ${\rho^{*}}$, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
4.1. KL divergence: no gradient dominance
We first present the following negative result concerning the gradient dominance condition for the KL-divergence.
@@ 4.1. KL divergence: no gradient dominance
(4.2) $\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-{\rm KL}[\rho\|{\rho^{*}}]\Bigr{)}^{2}\mathrm{d}\theta\leq\epsilon\rm{KL}[\rho\|{\rho^{*}}].$
Moreover, given any $\epsilon^{\prime}>0$, we can also construct $\rho$ in the neighborhood of ${\rho^{*}}$ defined as KL$[\rho\|{\rho^{*}}]<\epsilon^{\prime}$, such that (4.2) holds.
Proof.
We denote
(4.3) $\mathcal{G}(\rho,{\rho^{*}})=\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-{\rm KL}[\rho\|{\rho^{*}}]\Bigr{)}^{2}\mathrm{d}\theta.$
We first consider 1D Gaussian densities, $\rho(\theta)=\mathcal{N}(\theta;\mu,\sigma)$ and ${\rho^{*}}(\theta)=\mathcal{N}(\theta;0,1)$. We can calculate both sides of (4.2):
@@
$\frac{\mathcal{G}(\rho,{\rho^{*}})}{{\rm KL}[\rho\|{\rho^{*}}]}=\frac{\frac{1}{2}\Bigl{(}\frac{1}{M^{2}}-1\Bigr{)}^{2}+1}{\frac{1}{2M^{2}}+\log M-\frac{1}{2}+\frac{M^{2}}{2}}\leq\frac{3}{M^{2}}\leq\epsilon.$
Here we used the fact that $\frac{1}{2M^{2}}+\log M\geq\frac{1}{2}$ when $M\geq 1$.
Then we consider the general case. Given any ${\rho^{*}}$, we construct
(4.4) $\rho(\theta)={\rho^{*}}(\theta)(x_{1}\mathbbm{1}_{S^{c}_{r}}(\theta)+x_{2}\mathbbm{1}_{S_{r}}(\theta))$
where $\mathbbm{1}_{A}$ is the indicator function of a set $A$, and $S_{r}$ denotes the sphere with radius $r$. Then, similar to the proof of Theorem 3.3, we have the flexibility to select any $x_{1}>1>x_{2}>0$; in particular, we can adjust $r$ to ensure $\rho$ is a density function.
We calculate both sides of eq. 4.2 as follows:
@@
[…]
(4.5a) ${\rm KL}[\rho\|{\rho^{*}}]\leq$ $\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}\leq\log x_{1}=\epsilon^{\prime}$
(4.5b) $\frac{\mathcal{G}(\rho,{\rho^{*}})}{{\rm KL}[\rho\|{\rho^{*}}]}=$ $\frac{(1-\frac{1}{x_{1}})(\frac{1}{x_{2}}-1)\bigl{(}\log\frac{x_{1}}{x_{2}}\bigr{)}^{2}}{(\frac{1}{x_{1}}-\frac{1}{x_{2}})\bigl{(}(\frac{1}{x_{2}}-1)\log x_{1}+(1-\frac{1}{x_{1}})\log x_{2}\bigr{)}}$
@@
$<$ $\frac{(M+\epsilon^{\prime})^{2}}{(e^{M}-1)\epsilon^{\prime}-(1-e^{-\epsilon^{\prime}})M}<\frac{(2M)^{2}}{\epsilon^{\prime}e^{M}-2M}\epsilon.$
In the last inequality, we used that when $M\geq 20$, we have $e^{M/2}\geq 6M^{2}$, which implies
$4M^{2}\leq\epsilon\epsilon^{\prime}e^{M/2}6M^{2}-2M^{2}\leq(\epsilon e^{M}-2M)\epsilon^{\prime}.$
Finally, using the same mollifying approach in the proof of Theorem 3.3, we can smooth $\rho$ in (4.4) and obtain a anologous result. The proof is complete.
4.2. Sufficient and necessary conditions for gradient dominance
Moreover, we have the following sufficient and necessary condition on $f$ to satisfy the gradient dominance in Theorem 4.2. Since that both sides of the gradient dominance inequality (4.1) remain the same under the transformation $f(x)\rightarrow f(x)+c(x-1)$ for any $c$, we can further assume $f^{\prime}(1)=0$.

### v2

@@ 3.3 Geodesic convexity in the Fisher-Rao geometry
Therefore, the Hessian is not positive definite. Finally, we can smooth $\rho$ in (47) by introducing the mollifier:
@@
which has a compact support within $B_{\delta}$ and satisfies $\int\phi_{\delta}(\theta)\mathrm{d}\theta=1$. The smoothed density is constructed using convolution, through
[…]
Here, $x_{3}$ is a scalar determined by $\int\rho_{\delta}(\theta)\mathrm{d}\theta=1$. Note that $x_{1}>1>x_{2}>0$. It holds that $x_{2}\leq x_{3}\leq x_{1}$ and $x_{2}\leq\frac{\rho_{\delta}}{{\rho^{*}}}\leq x_{1}$. Since $\rho_{\delta}(\theta)$ only deviates from $\rho$ in (47) within the compact band $B_{r+2\delta}\backslash B_{r-2\delta}$, the last inequality in Eq. 49 remains valid when $\delta$ is sufficiently small.
Theorem 3.1 demonstrates that the KL divergence, a specific instance of an $f$-divergence, is not geodesically convex. This finding implies that $f$-divergences generally do not possess this property. Consequently, we cannot rely on convexity to establish exponential convergence for Fisher-Rao gradient flows of the KL divergence. In Section 5, we will introduce alternative conditions that enable us to prove convergence results.
Conversely, we will now establish a sufficient condition on $f$ that ensures the geodesic convexity of its corresponding $f$-divergence.
@@
Assume $f\in C^{2}(0,+\infty)$ is convex, satisfies $f(1)=0$ and $xf^{\prime}(x)$ is concave. Then, for any smooth, positive target density $\rho^{*}$, the $f$-divergence is geodesically convex within the Fisher-Rao geometry. Furthermore, when $f^{\prime\prime}(1)>0$, such $f$-divergence is $\alpha_{f}$-strongly geodesically convex for any $\rho^{*}$}, where $\alpha_{f}>0$ is a constant that depends solely on $f$. These assumptions are satisfied for $f^{\prime\prime}(x)=x^{p}$ with $p\leq-2$.
Proof 3.4.
@@
(50) $\mathbb{E}_{\rho}[f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)]=\mathbb{E}_{{\rho^{*}}}\Bigl[\frac{\rho}{{\rho^{*}}}f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\Bigr]\leq\mathbb{E}_{{\rho^{*}}}\bigl[\frac{\rho}{{\rho^{*}}}\bigr]f^{\prime}\Bigl(\mathbb{E}_{{\rho^{*}}}\bigl[\frac{\rho}{{\rho^{*}}}\bigr]\Bigr)=f^{\prime}(1).$
We recall the expression of Hessian operator from (42b)
(51) $\begin{split}&\phantom{{}={}}g_{\rho}^{\rm FR}\bigl(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma\bigr)\\ &=\frac{1}{2}\int\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}\Bigl(2\frac{\rho}{{\rho^{*}}}f^{\prime\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)+f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)-\mathbb{E}_{\rho}f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\Bigr)\mathrm{d}\theta\\ &\geq\frac{1}{2}\int\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}\Bigl(2\frac{\rho}{{\rho^{*}}}f^{\prime\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)+f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)-f^{\prime}(1)\Bigr)\mathrm{d}\theta,\end{split}$
where in the last line we used (50). We introduce the function
(52) $h(x)=2xf^{\prime\prime}(x)+f^{\prime}(x)-f^{\prime}(1)=2\bigl(xf^{\prime}(x)\bigr)^{\prime}-f^{\prime}(x)-f^{\prime}(1).$
@@
(53) $g_{\rho}^{\rm FR}\bigl(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma\bigr)\geq\frac{1}{2}\int\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}h\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\mathrm{d}\theta$
Due to our assumptions, $xf^{\prime}(x)$ and f(x)-f(x) are both concave, so $(xf^{\prime}(x))^{\prime}$ and $f^{\prime}(x)$ are non-increasing functions of $x$. It follows that $h(x)$ in (52) is non-increasing. Therefore, when $x\leq 1$, we have $h(x)\geq h(1)=2f^{\prime\prime}(1)\geq 0,$ and when $x\geq 1$, we have $h(x)\geq f^{\prime}(x)-f^{\prime}(1)\geq 0.$ Thus, $h(x)\geq 0$. This together with (53) implies that
$g_{\rho}^{\rm FR}\bigl(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma\bigr)\geq 0,$
for any tangent vector $\sigma$. This shows that the $f$-divergence is geodesically convex.
Next we prove that when $f^{\prime\prime}(1)>0$, the $f$-divergence is $\alpha_{f}$-strongly geodesically convex. To do so, we need a stronger estimate of the function $h(x)$. When $f^{\prime\prime}(1)>0$, there exists $\delta_{f}>0$, such that for $x\in[1,1+\delta_{f}]$, $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x)$. This inequality, combined with the preceding discussion, implies that
(54) $h(x)\geq\begin{cases}2f^{\prime\prime}(1)&0<x\leq 1\\ 2xf^{\prime\prime}(x)\geq f^{\prime\prime}(1)&1<x\leq 1+\delta_{f}\\ f^{\prime}(x)-f^{\prime}(1)=\int_{1}^{x}f^{\prime\prime}(\xi){\rm d}\xi\geq\delta_{f}\frac{f^{\prime\prime}(1)}{2}&x\geq 1+\delta_{f}\end{cases}.$
Consequently, $h(x)\geq\alpha_{f}>0$, where $\alpha_{f}=\min\{f^{\prime\prime}(1),\delta_{f}\frac{f^{\prime\prime}(1)}{2}\}$ is a constant that depends only on $f$. Thus, when $f^{\prime\prime}(1)>0$, combining (54) and (53), we conclude that the $f$-divergence is $\alpha_{f}$-strongly geodesically convex, i.e.,
@@
Finally, for $f^{\prime\prime}(x)=x^{p}$ with $p\leq-2$, $f$ is convex on $(0,+\infty)$. Moreover, $xf^{\prime}(x)$ is concave, since $(xf^{\prime}(x))^{\prime\prime}=2f^{\prime\prime}(x)+xf^{\prime\prime\prime}(x)=(2+p)x^{p}\leq 0$ for $x\in(0,+\infty)$. Thus, all assumptions are satisfied for $f$ with $f^{\prime\prime}(x)=x^{p}$ where $p\leq-2$.
4 Functional Inequality: Gradient Dominance Condition
In this section, we study the gradient dominance condition (11) of the $f$-divergences in the Fisher-Rao geometry. Counterexamples will be presented to illustrate that the KL-divergence does not satisfy the gradient dominance condition. Subsequently, we provide necessary and sufficient conditions to identify the functions $f$ that ensure the gradient dominance condition is met.
@@ 4 Functional Inequality: Gradient Dominance Condition
The gradient dominance condition can be formulated as follows:
@@
The function $f$ is said to satisfy the gradient dominance condition if there exists a constant $\alpha_{f}>0$, depending only on $f$, such that
(55) $\int\rho\left(f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)-\mathbb{E}_{\rho}\Bigl[f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\Bigr]\right)^{2}\mathrm{d}\theta\geq\alpha_{f}D_{f}[\rho\|{\rho^{*}}],$
for any smooth positive densities $\rho$ and $\rho^{*}$}, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
4.1 KL divergence: no gradient dominance
We first present the following negative result concerning the gradient dominance condition for the KL divergence. In this case, $f(x)=x\log x$ and $f^{\prime}(x)=\log x+1$.
@@ 4.1 KL divergence: no gradient dominance
(56) $\int\rho\Bigl(\log\frac{\rho}{{\rho^{*}}}-{\rm KL}[\rho\|{\rho^{*}}]\Bigr)^{2}\mathrm{d}\theta\leq\epsilon\rm{KL}[\rho\|{\rho^{*}}].$
[…]
Proof 4.2.
We denote $\mathcal{G}(\rho,{\rho^{*}})=\int\rho\Bigl(\log\frac{\rho}{{\rho^{*}}}-{\rm KL}[\rho\|{\rho^{*}}]\Bigr)^{2}\mathrm{d}\theta.$ We first consider simple 1D Gaussian densities as a pedagogical example: $\rho(\theta)=\mathcal{N}(\theta;\mu,\sigma)$ and ${\rho^{*}}(\theta)=\mathcal{N}(\theta;0,1)$. We can calculate both sides of (56) similar to Eq. 45:
@@
$\frac{\mathcal{G}(\rho,{\rho^{*}})}{{\rm KL}[\rho\|{\rho^{*}}]}=\frac{\frac{1}{2}\Bigl(\frac{1}{M^{2}}-1\Bigr)^{2}+1}{\frac{1}{2M^{2}}+\log M-\frac{1}{2}+\frac{M^{2}}{2}}\leq\frac{3}{M^{2}}\leq\epsilon,$
and thus, the gradient dominance condition is violated. Here we used the fact that $\frac{1}{2M^{2}}+\log M\geq\frac{1}{2}$ when $M\geq 1$.
Then, we consider the general case. Given any $\rho^{*}$}, we construct
(57) $\rho(\theta)={\rho^{*}}(\theta)\bigl(x_{1}\mathbbm{1}_{B^{c}_{r}}(\theta)+x_{2}\mathbbm{1}_{B_{r}}(\theta)\bigr).$
Then, similar to the proof of Theorem 3.1, we have the flexibility to select any $x_{1}>1>x_{2}>0$; in particular, we can adjust $r$ to ensure $\rho$ is a density function.
We calculate both sides of Eq. 56 as follows:
@@
If we choose $x_{1}=e^{\epsilon^{\prime}}$ with $\epsilon^{\prime}>0$ and $x_{2}=e^{-M}$ with $M>\max\{\epsilon,\,\epsilon^{\prime},\,2\log\frac{1}{\epsilon\epsilon^{\prime}},\,20\}$, we have
(58a) ${\rm KL}[\rho\|{\rho^{*}}]\leq$ $\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}\leq\log x_{1}=\epsilon^{\prime}$
(58b) $\frac{\mathcal{G}(\rho,{\rho^{*}})}{{\rm KL}[\rho\|{\rho^{*}}]}=$ $\frac{(1-\frac{1}{x_{1}})(\frac{1}{x_{2}}-1)\bigl(\log\frac{x_{1}}{x_{2}}\bigr)^{2}}{(\frac{1}{x_{2}}-\frac{1}{x_{1}})\bigl((\frac{1}{x_{2}}-1)\log x_{1}+(1-\frac{1}{x_{1}})\log x_{2}\bigr)}$
@@
$<$ $\frac{(M+\epsilon^{\prime})^{2}}{(e^{M}-1)\epsilon^{\prime}-(1-e^{-\epsilon^{\prime}})M}<\frac{4M^{2}}{\epsilon^{\prime}e^{M}-2M}\epsilon.$
In the last inequality, we used that when $M\geq 20$, $e^{M/2}\geq 6M^{2}$, and $4M^{2}\leq\epsilon\epsilon^{\prime}e^{M/2}6M^{2}-2M^{2}\leq(\epsilon e^{M}-2M)\epsilon^{\prime}$.
Lastly, by applying the same mollifier approach used in the proof of Theorem 3.1, we can smooth $\rho$ in (57) to obtain an analogous result. This completes the proof.
also note that [lu2019accelerating, Theorem 3.3][lu2022birth, Theorem 2.4] provide a sufficient condition for a density $\rho$ to satisfy the gradient dominance condition, namely that $\inf_{\theta}\frac{\rho(\theta)}{{\rho^{*}}(\theta)}>0$, i.e., that the ratio has a positive lower bound. Our counterexamples here do not satisfy this condition. Specifically, for the sequence of densities $\rho$ parameterized by $M$, the constant for the gradient dominance condition worsens and tends to infinity as $M\rightarrow\infty$. This blow-up can still occur even when $\rho$ remains within an $\epsilon^{\prime}$-neighborhood of $\rho^{*}$.
4.2 Sufficient and necessary conditions for gradient dominance
In this section, we present a sufficient and necessary condition on $f$ so that the gradient dominance is satisfied.
Since both sides of the gradient dominance condition (55) remain the same under the transformation $f(x)\rightarrow f(x)+c(x-1)$ for any $c$, we can further assume $f^{\prime}(1)=0$.